A steel wire in a piano has a length of 0.560 m and a mass of 4.900 x 10-3 kg. To what tension must this middle C(c-261.6 Hz on the chromatic musikal scale)? were be wetched so tha se hramental etesten estepona, to
Answer #1

By using,

f = (1/2L)*sqrt(F/(m/L))

F = (f*2L)^2 *(m/L)

= 4f^2 mL

= 4*(261.6)^2 *4.9*10^-3*0.56

= 751.14 N